# Composing Codensity Bisimulations

### v1

@@ A.3.3 Bisimulation Metric
Therefore it’s enough to show that two inequalities
$(lhs)\geq\sup_{(a_{1},b_{1})\in\lambda^{\mathcal{P}}(A_{1},B_{1})}\inf_{(a_{2},b_{2})\in\lambda^{\mathcal{P}}(A_{2},B_{2})}[T,\sigma]^{\mathbf{\Omega}}(P,Q)((a_{1},b_{1}),(a_{2},b_{2}))$
and
@@ A.4 Omitted Proofs for Section 8
$\lambda^{*}[F,\tau]^{\mathbf{\Omega}}[T,\sigma]^{\mathbf{\Omega}}$ $=\lambda^{*}R^{p,\mathbf{\Omega}}\circ\mathrm{Sp}^{A}(F,\tau)\circ\mathrm{Sp}^{A}(T,\sigma)\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$
(by the condition (2) in Theorem 6.2)
$=R^{p,\mathbf{\Omega}}\circ\lambda^{*}\mathrm{Sp}^{A}(F,\tau)\circ\mathrm{Sp}^{A}(T,\sigma)\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$
(since $R^{p,\mathbf{\Omega}}$ preserves cartesian morphisms)
$=R^{p,\mathbf{\Omega}}\circ\mathrm{Sp}^{A}(T,\sigma)\circ\mathrm{Sp}^{A}(F^{N},\tau^{N})\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$
(by the condition (1) in Theorem 6.2 and discussion after Theorem 6.2)
@@
$\iff\textstyle\bigsqcup_{\langle P_{i}\rangle_{i\in N}\in\mathcal{V}}[T,\sigma]^{\mathbf{\Omega}}(\langle P_{i}\rangle_{i\in N})$
$\sqsubseteq(T(\langle c_{i}\rangle_{i\in N}))^{*}R^{p,\mathbf{\Omega}}\circ\mathrm{Sp}^{A}(T\circ F^{N},\sigma\mathbin{\bullet}(T\circ\tau^{N}))\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}(\langle\bigsqcup_{\langle P_{i}\rangle_{i\in N}\in\mathcal{V}}P_{i}\rangle_{i\in N})$
$=\bigsqcap_{a\in A,\langle k_{i}\in\mathbb{E}(\bigsqcup_{\langle P_{i}\rangle_{i}\in\mathcal{V}}P_{i},\mathbf{\Omega})\rangle_{i\in N}}\big{(}\sigma\circ T(\langle\tau_{a}\circ Fpk_{i}\circ c_{i}\rangle_{i\in N})\big{)}^{*}\mathbf{\Omega}$
$\iff\forall\langle P_{i}\rangle_{i\in N}\in\mathcal{V}.~{}\forall a\in A.~{}\forall\langle k_{i}\colon\bigsqcup_{\langle P_{i}^{\prime}\rangle_{i}\in\mathcal{V}}P_{i}^{\prime}\dot{\rightarrow}\mathbf{\Omega}\rangle_{i\in N}.~{}[T,\sigma]^{\mathbf{\Omega}}(\langle P_{i}\rangle_{i\in N})\sqsubseteq\big{(}\sigma\circ T(\langle\tau_{a}\circ Fk_{i}\circ c_{i}\rangle_{i\in N})\big{)}^{*}\mathbf{\Omega}$
$\iff\forall\langle P_{i}\rangle_{i\in N}\in\mathcal{V}.~{}\forall a\in A.~{}\forall\langle k_{i}\in\mathbb{C}(X_{i},\Omega)\rangle_{i\in N}.$
$\phantom{\iff}\bigl{(}\forall\langle P_{i}^{\prime}\rangle_{i\in N}\in\mathcal{V}.~{}\forall i\in N~{}k_{i}\colon P_{i}^{\prime}\dot{\rightarrow}\mathbf{\Omega}\Rightarrow[T,\sigma]^{\mathbf{\Omega}}(\langle P_{i}\rangle_{i\in N})\sqsubseteq\big{(}\sigma\circ T(\langle\tau_{a}\circ Fk_{i}\circ c_{i}\rangle_{i\in N})\big{)}^{*}\mathbf{\Omega}\bigr{)}$
$\iff\forall\langle P_{i}\rangle_{i\in N}\in\mathcal{V}.~{}\forall a\in A.~{}\forall\langle k_{i}\in\mathbb{C}(X_{i},\Omega)\rangle_{i\in N}.$
$\phantom{\iff}\bigl{(}[T,\sigma]^{\mathbf{\Omega}}(\langle P_{i}\rangle_{i\in N})\not\sqsubseteq\big{(}\sigma\circ T(\langle\tau_{a}\circ Fk_{i}\circ c_{i}\rangle_{i\in N})\big{)}^{*}\mathbf{\Omega}\Rightarrow\exists\langle P_{i}^{\prime}\rangle_{i\in N}\in\mathcal{V}.~{}\exists i\in N~{}k_{i}\colon P_{i}^{\prime}\dot{\not\rightarrow}\mathbf{\Omega}\bigr{)}.$
Proof of Theorem 8.9.

### v2

@@ A.3.3 Bisimulation Metric
Therefore it’s enough to show that two inequalities:
  • (lhs) $\geq\sup_{(a_{1},b_{1})\in\lambda^{\mathcal{P}}(A_{1},B_{1})}\inf_{(a_{2},b_{2})\in\lambda^{\mathcal{P}}(A_{2},B_{2})}[T,\sigma]^{\mathbf{\Omega}}(P,Q)((a_{1},b_{1}),(a_{2},b_{2}))$,
@@ A.4 Omitted Proofs for Section 8
$\lambda^{*}[F,\tau]^{\mathbf{\Omega}}[T,\sigma]^{\mathbf{\Omega}}$
$=\lambda^{*}R^{p,\mathbf{\Omega}}\circ\mathrm{Sp}^{A}(F,\tau)\circ\mathrm{Sp}^{A}(T,\sigma)\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$ by the condition (2) in Theorem 6.2
$=R^{p,\mathbf{\Omega}}\circ\lambda^{*}\mathrm{Sp}^{A}(F,\tau)\circ\mathrm{Sp}^{A}(T,\sigma)\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$ since $R^{p,\mathbf{\Omega}}$ preserves cartesian morphisms
$=R^{p,\mathbf{\Omega}}\circ\mathrm{Sp}^{A}(T,\sigma)\circ\mathrm{Sp}^{A}(F^{N},\tau^{N})\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$ by the condition (1) in Theorem 6.2,
and discussion after Theorem 6.2
@@
if and only if
$\bigsqcup_{\langle P_{i}\rangle_{i\in N}\in\mathcal{V}}[T,\sigma]^{\mathbf{\Omega}}(\langle P_{i}\rangle_{i\in N})\sqsubseteq(T(\langle c_{i}\rangle_{i\in N}))^{*}R^{p,\mathbf{\Omega}}\circ\mathrm{Sp}^{A}(T\circ F^{N},\sigma\mathbin{\bullet}(T\circ\tau^{N}))\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}(\langle\bigsqcup_{\langle P_{i}\rangle_{i\in N}\in\mathcal{V}}P_{i}\rangle_{i\in N})$
[…]
if and only if
$\forall\langle P_{i}\rangle_{i\in N}\in\mathcal{V}.~{}\forall a\in A.~{}\forall\langle k_{i}\colon\bigsqcup_{\langle P_{i}^{\prime}\rangle_{i}\in\mathcal{V}}P_{i}^{\prime}\dot{\rightarrow}\mathbf{\Omega}\rangle_{i\in N}.~{}[T,\sigma]^{\mathbf{\Omega}}(\langle P_{i}\rangle_{i\in N})\sqsubseteq\big{(}\sigma\circ T(\langle\tau_{a}\circ Fk_{i}\circ c_{i}\rangle_{i\in N})\big{)}^{*}\mathbf{\Omega}$
$\text{if and only if }\forall\langle P_{i}\rangle_{i\in N}\in\mathcal{V}.~{}\forall a\in A.~{}\forall\langle k_{i}\in\mathbb{C}(X_{i},\Omega)\rangle_{i\in N}.$
[…]
$\text{if and only if }\forall\langle P_{i}\rangle_{i\in N}\in\mathcal{V}.~{}\forall a\in A.~{}\forall\langle k_{i}\in\mathbb{C}(X_{i},\Omega)\rangle_{i\in N}.$
[…]
Lemma A.10.
Let $\mathcal{V}\subseteq\prod_{i\in N}\mathrm{Obj}(\mathbb{E}_{X_{i}})$ be a set. If a set $\mathcal{V}$ is an invariant of a codensity game $\mathcal{G}_{c}$ for $\mathbf{D}$ then so is $\overline{\mathcal{V}}$.
Proof.
Suppose $\mathcal{V}$ is an invariant of $\mathcal{G}_{c}$ for $\mathbf{D}$. To show that $\overline{\mathcal{V}}$ is an invariant of $\mathcal{G}_{c}$ for $\mathbf{D}$, we need to check: for each $(a\in A,k\colon X\to\Omega(a))$ such that $\tau_{a}\circ Fk\circ c\colon\bigsqcup_{P\in\mathcal{V}}P\dot{\not{\rightarrow}}\mathbf{\Omega}(a)$, there exists $P^{\prime}\in\mathcal{V}$ such that $k\colon P^{\prime}\dot{\not{\rightarrow}}\Omega(a)$. It is true because $\tau_{a}\circ Fk\circ c\colon\bigsqcup_{P\in\mathcal{V}}P\dot{\not{\rightarrow}}\mathbf{\Omega}(a)$ implies $\exists~{}P^{\prime}\in\mathcal{V}.~{}\tau_{a}\circ Fk\circ c\colon P^{\prime}\dot{\not{\rightarrow}}\mathbf{\Omega}(a)$ and $\mathcal{V}$ is an invariant. ∎